# A Data-Centric Revisit of Pre-Trained Vision Models for Robot Learning

### v1

@@ 1 Introduction
Pre-trained vision models (PVMs) have become fundamental building blocks in modern computer vision and robotics. While these models have demonstrated remarkable success in traditional vision tasks, their optimal application in robot learning remains an open challenge. Recent works [55, 45] have shown promising results using masked autoencoders (MAE) [29] pre-trained on ego-centric data (e.g., Ego4D [24]). However, two critical questions remain unexplored: 1) Is MAE the optimal pre-training method for robot learning? 2) Is ego-centric data the best choice for visuomotor pre-training? As suggested by [19], learning on traditional datasets like ImageNet [20] can obtain better representations. Additionally, scene-centric data (e.g., COCO [41] and Open Images [39]) are also relevant to robot contexts and more information-rich. Web-crawled data (e.g., CC12M [14]) offer another alternative, being easier to collect and more scalable.
@@
Motivated by these insights, we introduce SlotMIM (Fig. 5), a method that repurposes and integrates masked image modeling (MIM) and contrastive learning for effective representation learning from NOC datasets. The core idea of SlotMIM is to group patch-level image tokens into object-level feature abstractions, referred to as “slots”, thereby decomposing NOC data into object-centric slots so that object-centric techniques can be effectively applied. To make patch-level tokens more semantically aware for subsequent grouping, we enhance MIM with cross-view consistency regularization. Additionally, we introduce a semantic bottleneck, which reduces the number of prototypes to encourage the emergence of semantic and objectness at patch-level token representations (see Fig. 4). Building on these semantically enriched patch tokens, we apply attentive pooling over the learned patch-level features, using prototypes to initialize object representations, thereby grouping patches into object-level slots. Contrastive learning [18] is then applied to these slots to promote the discriminativeness of the learned representations.
Our scaled experiments (Fig. 6) extend the analysis to million-scale datasets, revealing several surprising findings about the relationship between pre-training data and model performance. First, we observe an unexpected inverse scaling trend: while MAE’s performance improves with more pre-training data, DINO and iBOT show degraded performance when scaled up. This indicates over-compression of representations: the learning objective reduces properties that are useful for manipulation tasks. Interestingly, we find that SlotMIM trained on ego-centric data learns to discover concepts adaptive to the pre-training data distribution, avoiding over-compression and consistently improving with more data. The resulting models—evaluated on six diverse tasks including Franka Kitchen, Meta-World, ObjectNav, ImageNav, COCO object detection and instance segmentation, and ADE20K semantic segmentation—demonstrate that SlotMIM consistently outperforms existing methods with better efficiency and scalability. Specifically, when pre-trained with just 241K samples, it already outperforms prior methods that used over 1M samples, including established approaches like MVP and VC-1. More significantly, while other methods plateau or degrade with increased NOC data, SlotMIM continues to improve and surpass previous methods that used 3$\times$ more data on both ADE20K and COCO benchmarks. This suggests that NOC data, when properly leveraged, can be a more scalable and efficient learning resource than previously thought.
@@ 2 Related Work
Several recent works have tackled self-supervised learning on non-object centric data [67, 22, 74, 68, 30, 31, 69, 3]. Among them, the majority focuses on learning from scene-centric datasets, and benefiting scene understanding tasks. To solve this problem, the idea is primarily either to build pretext tasks on dense featuremaps [68, 74, 3], object-centric groups [30, 31, 69], or specialized data augmentation techniques [22, 67]. In addition, some works have also explored learning from uncurated datasets [11, 66, 3, 52], which mainly correspond to large-scale web-crawled datasets, and a main topic is data deduplication [52]. Our work contributes to this line of research by systematically studying the performance of self-supervised learning on multiple types of NOC data.
@@ 3.2 Evaluation Setting
Figure 3: Behavior cloning with attentive probing. An additional token is trained with cross attention to gather information from all patch tokens from the backbone.
@@ 4.2 SlotMIM Framework
iBOT uses a set of prototype embeddings $\mathcal{C}=\{{\bm{c}}_{c}\}_{c=1}^{C}$ to cluster image patches into $C$ groups, assigning each patch token a soft one-hot encoding $p_{\theta}({\bm{x}}_{i})$ of its cluster membership. While iBOT typically uses $C=8192$ to capture fine patterns, we find a much smaller $C$ (e.g., 512 for COCO) better serves object discovery by creating an information bottleneck that encourages learning compositional object concepts. As shown in Fig. 4(a), iBOT’s clusters are very fine-grained (2nd row), but objectness emerges with less prototypes (3rd row). However, these patterns still lack semantic meaning and can fragment single objects. Additionally, matching discovered objects between views remains difficult as their semantics vary despite assigned to the same prototype (Fig. 4(b), left). Both issues suggest the need for semantic-level prototypes.
@@ 4.3 Ablation Study
We conduct an ablation study to analyze key components of SlotMIM. All models are trained on COCO+ for 800 epochs. As shown in Tab. 3, we evaluate each model variant using three metrics: $k$-NN accuracy on ImageNet classification, mean IoU on ADE20K semantic segmentation, and Jaccard index on VOC2012. We also report $\overline{K}$, the average number of objects/stuff regions discovered per image. The results reveal that Jaccard index positively correlates with representation quality, indicating that better objectness leads to stronger representations. Adding MIM improves downstream segmentation performance (comparing rows 1 and 2). The cross-view consistency loss and slot contrastive loss both contribute significantly to objectness (rows 2, 3, 5). Additionally, the within-view loss acts as an effective regularizer that further enhances the learned representations (rows 4, 5). We provide detailed ablation studies on hyperparameters in the appendix.
@@ Acknowledgements
This work has been supported by Hong Kong Research Grant Council—Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422), and RGC Matching Fund Scheme (RMGS). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.

### v2

@@ 1 Introduction
Pre-trained vision models (PVMs) have become fundamental building blocks in modern computer vision and robotics. While these models have demonstrated remarkable success in traditional vision tasks, their optimal application in robot learning remains an open challenge. Recent works [55, 45] have shown promising results using masked autoencoders (MAEs) [29] pre-trained on ego-centric data (e.g., Ego4D [24]). However, as suggested by [19], learning on (single-)object-centric datasets like ImageNet [20] can obtain better representations than on ego-centric data. Additionally, scene-centric data (e.g., COCO [41] and Open Images [39]) are also relevant to robot contexts and more information-rich. Web-crawled data (e.g., CC12M [14]) offer another alternative, being easier to collect and more scalable. Towards scaling up PVMs for real-world robotic applications, it is important to explore diverse training data sources and revisit learning algorithms accordingly. Our study is thus motivated by two questions: 1) Is MAE the optimal pre-training method for robot learning? 2) Is ego-centric data the best choice for visuomotor pre-training?
@@
Motivated by these insights, we introduce SlotMIM (Fig. 5), a method that repurposes and integrates masked image modeling (MIM) and contrastive learning for effective representation learning from NOC datasets. The core idea of SlotMIM is to group patch-level image tokens into object-level feature abstractions, referred to as “slots”, thereby decomposing NOC data into object-centric slots so that object-centric techniques can be effectively applied. To make patch-level tokens more semantically aware for subsequent grouping, we enhance MIM with cross-view consistency regularization regarding prototype assignments. Additionally, we introduce a semantic bottleneck, which reduces the number of prototypes to encourage the emergence of semantic and objectness at patch-level token representations (see Fig. 4). Building on these semantically enriched patch tokens, we apply attentive pooling over the learned patch-level features, using prototypes to initialize object representations, thereby grouping patches into object-level slots. Contrastive learning [18] is then applied to these slots to improve the discriminativeness of the learned representations. Together, these designs enable us to perform effective representation learning from NOC data.
Our scaled experiments (Fig. 6) extend the analysis to million-scale datasets, revealing several surprising findings about the relationship between pre-training data and model performance. First, we observe an unexpected inverse scaling trend: while MAE’s performance improves with more pre-training data, DINO and iBOT show degraded performance when scaled up. This indicates over-compression of representations: the learning objective reduces properties that are useful for manipulation tasks. Interestingly, we find that SlotMIM trained on ego-centric data learns to discover concepts adaptive to the pre-training data distribution, avoiding over-compression and consistently improving with more data. The resulting models—evaluated on six diverse tasks including Franka Kitchen, Meta-World, ObjectNav, ImageNav, COCO object detection and instance segmentation, and ADE20K semantic segmentation—demonstrate that SlotMIM consistently outperforms existing methods with better efficiency and scalability. Specifically, when pre-trained with just 241K samples, it already outperforms prior methods that used over 1M samples, including established approaches like MVP and VC-1. More significantly, while other methods saturate or degrade with increased NOC data, SlotMIM continues to improve and surpass previous methods that used 3$\times$ more data on both ADE20K and COCO benchmarks. This suggests that NOC data, when properly leveraged, can be a more scalable and efficient learning resource than previously thought, enabling new possibilities for scaling up PVMs.
@@ 2 Related Work
Several recent works have tackled self-supervised learning on non-object centric data [67, 22, 74, 68, 30, 31, 69, 3]. Among them, the majority focuses on learning from scene-centric data and benefiting scene understanding tasks. The idea is primarily either to build pretext tasks on dense feature maps [68, 74, 3], object-centric groups [30, 31, 69], or specialized data augmentation techniques [22, 67]. In addition, some works have also explored learning from uncurated datasets [11, 66, 3, 52], which mainly correspond to large-scale web-crawled datasets, and a main topic is data deduplication [52]. Our work contributes to this line of research by systematically studying the performance of self-supervised learning on multiple types of NOC data.
@@ 3.2 Evaluation Setting
Figure 3: Behavior cloning with attentive probing. An additional token is trained with cross-attention (trainable) to gather information from all patch tokens from the backbone (frozen), and fed to the policy to learn from expert demonstrations via behavior cloning.
@@ 4.2 SlotMIM Framework
iBOT uses a set of prototype embeddings $\mathcal{C}=\{{\bm{c}}_{c}\}_{c=1}^{C}$ to cluster image patches into $C$ groups, assigning each patch token a soft one-hot encoding $p_{\theta}({\bm{x}}_{i})$ of its cluster membership. While iBOT typically uses $C=8192$ to capture fine patterns, we find a much smaller $C$ (e.g., 512 for COCO) better serves object discovery by creating an information bottleneck that encourages learning compositional object concepts. As shown in Fig. 4(a), iBOT’s clusters are very fine-grained (2nd row), but objectness emerges with fewer prototypes (3rd row). However, these patterns still lack semantic meaning and can fragment single objects. Additionally, matching discovered objects between views remains difficult as their semantics vary despite being assigned to the same prototype (Fig. 4(b), left). Both issues suggest the need for semantic-level prototypes.
@@ 4.3 Ablation Study
We ablate on several key components of SlotMIM. All models are trained on COCO+ for 800 epochs. As shown in Tab. 3, we evaluate each model variant using three metrics: $k$-NN accuracy on ImageNet classification, mean IoU on ADE20K semantic segmentation, and Jaccard index on VOC2012. We also report $\overline{K}$, the average number of objects/stuff regions discovered per image. The results reveal that Jaccard index positively correlates with representation quality, indicating that better objectness leads to stronger representations. Adding MIM improves downstream segmentation performance (comparing rows 1 and 2). The cross-view consistency loss and slot contrastive loss both contribute significantly to objectness (rows 2, 3, 5). Additionally, the within-view loss acts as an effective regularizer that further enhances the learned representations (rows 4, 5). Additional ablation study are available in the appendix.
@@ Acknowledgements
This work has been supported in part by Hong Kong Research Grant Council — Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923), Theme-based Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.